# Möbius-Transformed Trapezoidal Rule

### v1

@@ Abstract.
This work was supported by the Research Council of Finland (decisions 338567, 348503, 353081 and 359181).
@@ 1. Introduction
Let us briefly review recent literature on numerical integration and function approximation for finitely smooth functions over unbounded domains. When the weight is a Gaussian probability density, the corresponding weighted Sobolev spaces have been considered by numerous authors; see, e.g., [7, 19, 6, 12, 13] and the references therein. Freud weights, one possible generalization of Gaussian weights, have also been studied in our general context [10, 9], as have unweighted Sobolev spaces with a certain decay condition [27]. All aforementioned articles use the same weight for defining the Sobolev space for the target function and measuring the error in the considered approximation, which is an assumption that can be dropped: [25] employs unweighted Sobolev spaces for the target function but presents weighted $L^{2}$ and $L^{\infty}$ error estimates, and [23] studies the relation between the two weights in a general framework. Table 1 summarizes the settings of those aforementioned papers whose foci are close to ours.
@@
The rest of this paper is organized as follows. In Section 2, we introduce and prove necessary concepts and useful lemmas related to monotonic Schwartz weights, Möbius transformations, and Sobolev spaces. Section 3 presents our main result for numerical integration, i.e., that the Möbius-transformed trapezoidal rule achieves the optimal rate of convergence. Section 4 extends this result to a randomized setting. In Section 5, we consider a problem of function approximation and prove the optimality of an algorithm that is based on combining the Möbius-transformed trapezoidal rule with trigonometric interpolation. Section 6 briefly considers a multidimensional extension of our method. Finally, Section 7 presents the concluding remarks.
@@ 2.1. Monotonic Schwartz weights
$x^{\alpha}\frac{|\omega^{(\beta)}(x)|^{r}}{\omega(x)}\leq(C^{\prime\prime\prime}_{\beta,m,\omega})^{r}x^{\alpha}\omega(x)^{(mr/(m+\beta)-1)}\longrightarrow 0$
@@
As (2.7) immediately follows from the definition of $\mathcal{S}$ for $\beta=0$, we may assume that $\beta\in\mathbb{N}$. A straightforward induction argument shows that for any $s>0$, the derivative $(\omega^{s})^{(\beta)}$ is a finite linear combination of terms of the form
@@ 2.3. Sobolev spaces
for some $1<q<\infty$ and $\alpha\in\mathbb{N}$, with $f^{(\tau)}$ denoting the $\tau$th weak derivative of $f$. The fact that $W^{\alpha,q}_{\rho}(\mathbb{R})$ is a Banach space (or a Hilbert space for $q=2$) for $\rho\in\mathcal{S}_{+}^{\rm mon}$ is a consequence of [21, Theorem 1.1 & Remark 4.10]. In what follows, we always consider the continuous representative of any given $f\in W^{\alpha,q}_{\rho}(\mathbb{R})$, which is possible for $\alpha\in\mathbb{N}$ due to the Sobolev embedding theorem and the inclusion $W^{\alpha,q}_{\rho}(\mathbb{R})\subset W^{1,q}_{\rm loc}(\mathbb{R})$. In Section 3, we specifically consider the case $q=2$ for numerical integration, even though the results also hold for any $q\in(1,\infty)$.
@@
The point evaluations at 00 and $2\pi$ in (2.3) are well-defined due to the trace theorem [1, Theorem 7.53] or the continuous embedding $W^{\alpha,q}(0,2\pi)\hookrightarrow C^{\alpha-1,\lambda}([0,2\pi])$ for Hölder indices $0<\lambda\leq 1-1/q$ [1, Part II in Theorem 5.4], demonstrating also that $W^{\alpha,q}_{\rm per}(0,2\pi)$ is a Banach space (or a Hilbert space for a $q=2$) as a closed subspace of $W^{\alpha,q}(0,2\pi)$. The closure of the smooth compactly supported test functions $C^{\infty}_{c}(0,2\pi)$ in the topology of $W^{\alpha,q}(0,2\pi)$ is denoted by $\mathring{W}^{\alpha,q}(0,2\pi)\subset W^{\alpha,q}_{\rm per}(0,2\pi)$. Note that in (2.14) the weak derivative $f^{(\tau)}$ is defined via partial integration of periodic functions on $\mathbb{T}$ (see, e.g., [29, Section 5.2]), whereas in (2.3), $f^{(\tau)}$ denotes the standard weak derivative on $(0,2\pi)\subset\mathbb{R}$, with $C^{\infty}_{c}(0,2\pi)$ serving as the space of test functions.
@@
For $1<q<\infty$ and $\alpha\in\mathbb{N}$, the spaces $W^{\alpha,q}(\mathbb{T})$ and $W^{\alpha,q}_{\rm per}(0,2\pi)$ can be identified via a bijective linear isometry $L:W^{\alpha,q}(\mathbb{T})\to W^{\alpha,q}_{\rm per}(0,2\pi)$.
@@
Let $\varphi:\mathbb{R}\ni\theta\mapsto(\cos\theta,\sin\theta)$ be the standard $2\pi$-periodic parametrization with respect to a polar angle $\theta$ for the embedding of $\mathbb{T}$ into $\mathbb{R}^{2}$. As $|\varphi^{\prime}|\equiv 1$, it is straightforward to conclude that the linear mapping $L:g\mapsto g\circ\varphi$ defines a bijective isometry from $W^{\alpha,q}(\mathbb{T})$ to
@@
for all $2\pi$-periodic $\psi\in C^{\infty}(\mathbb{R})$ (cf. [29, (5.5), (5.13) & Exercise 5.3.2]). Hence, the assertion follows if we prove that $\widetilde{W}^{\alpha,q}_{\rm per}(0,2\pi)=W^{\alpha,q}_{\rm per}(0,2\pi)$. This boils down to showing that the weak derivatives of $f\in W^{\alpha,q}_{\rm per}(0,2\pi)$ satisfy (2.16) and that those of $f\in\widetilde{W}^{\alpha,q}_{\rm per}(0,2\pi)$ are compatible with the conditions on point evaluations at 00 and $2\pi$ in (2.3).
@@
for all all $2\pi$-periodic $\psi\in C^{\infty}(\mathbb{R})$, which could be proved by, e.g., approximating $f$ with smooth functions, integrating by parts and employing the continuous embedding $W^{\alpha,q}(0,2\pi)\hookrightarrow C^{\alpha-1,\lambda}([0,2\pi])$ for $0<\lambda\leq 1-1/q$ (cf. [1, Theorem 3.18 & Part II of Theorem 5.4]). The claim now follows by choosing in turns $f\in W^{\alpha,q}_{\rm per}(0,2\pi)$ and $f\in\widetilde{W}^{\alpha,q}_{\rm per}(0,2\pi)$ in (2.3) and comparing with (2.3) and (2.16). ∎
@@ 3. Möbius-Transformed Trapezoidal Rule
For $\rho\in\mathcal{S}_{+}^{\rm mon}$, it holds that
(3.2) $\big{|}I_{\rho}(f)-Q_{\rho,n}(f)\big{|}\leq C_{\rho,\alpha}n^{-\alpha}\|f\|_{W^{\alpha,2}_{\rho}(\mathbb{R})},\quad\alpha\in\mathbb{N},$
@@
The proof relies on showing that $g=((f\rho)\circ\phi)\phi^{\prime}$ belongs to the $2\pi$-periodic $L^{2}$-based Sobolev space $W^{\alpha,2}_{\rm per}(0,2\pi)\cong W^{\alpha,2}(\mathbb{T})$ if $f\in W^{\alpha,2}_{\rho}(\mathbb{R})$ for $\alpha\in\mathbb{N}$. Indeed, after proving this, a classical result for the trapezoidal rule on periodic Sobolev spaces guarantees the convergence rate of order $n^{-\alpha}$; see, e.g.,  [2, Proposition 7.5.6] and [42, Theorem 2.4.1].
@@
Let us start with two auxiliary results that can be proved by straightforward induction arguments. First, the weak derivative $g^{(\tau)}$, $\tau\in\mathbb{N}_{0}$, on the open interval $(0,2\pi)$ is a finite linear combination of terms of the form
@@
Figure 2. Absolute integration error for the logistic weight and $f(x)=|x|^{p}$, which corresponds to $I_{\rho}(f)=-2\,p!\,{\rm Li}_{p}(-1)$ where ${\rm Li}$ denotes the polylogarithm [18, Corollary 4.2]. The blue line shows the error for the Gauss–Logistic quadrature from [11]. The Möbius-transformed trapezoidal rule (green) exhibits much faster convergence.
@@ 4. Randomized trapezoidal rule
The convergence rate $n^{-\alpha-1/2}$ of the RMSE for the randomized quadrature rule (4.1) is a direct consequence of Lemma 3.2 combined with the classical result by Bakhvalov, [3] or [20, Theorem 11] where multidimensional settings are considered. Since our setting is one-dimensional, we refer to [13, p. 1670] for a better bound on the RMSE for $2T$-periodic functions. For completeness, we formally state this result as follows:
@@ 5. $L^{p}_{\rho}$ approximation
is as small as possible for $0\leq p<q<\infty$.
@@
To deduce the convergence rate of Theorem 5.2, we generalize Lemma 3.2 to show that the transformed target function $g_{p}$ belongs to the periodic $L^{q}$-based Sobolev space $W^{\alpha,q}_{\rm per}(0,2\pi)\cong W^{\alpha,q}(\mathbb{T})$ for $q>p$. The proof of Theorem 5.2 then follows from a trigonometric interpolation result by Temlyakov [42, Theorem 2.7].
@@
A straightforward induction argument demonstrates that the weak derivative $g_{p}^{(\tau)}$, $\tau\in\mathbb{N}$, on the open interval $(0,2\pi)$ is a finite linear combination of terms of the form
@@
We follow the line of reasoning in [6, Proof of Theorem 3.3]. To this end, let $f\in W^{\alpha,q}_{\rm per}(0,1)$ and denote its $1$-periodic extension onto the real line by the same symbol. Since the extension $f$ has $\alpha-1$ continuous weak derivatives, it belongs to $W^{\alpha,q}_{\rm loc}(\mathbb{R})$ (cf. Proposition 2.3). For any $N\in\mathbb{N}$,
@@
The claim then follows from a lower bound for the approximation of periodic functions [42, Theorem 2.1.1], presented originally in [24]. ∎
@@ 6. Multivariate extension by componentwise transforms
Let $\alpha\in\mathbb{N}$, $\rho(\boldsymbol{x})=\prod_{k=1}^{d}\rho_{j}(x_{j})$ with $\rho_{j}\in\mathcal{S}_{+}^{\rm mon}$, $f\in W^{\alpha,2}_{\rho,\otimes}(\mathbb{R}^{d})$, and
@@
As the modified integrand function in (6.1) is in the periodic Sobolev space $W^{\alpha,2}_{\otimes}(\mathbb{T}^{d})$ that is norm-equivalent to the Sobolev space of dominating mixed smoothness $W^{\alpha,2}_{\rm mix}(\mathbb{T}^{d})$, one can employ in (6.1) good rank-$1$ lattice points [26, Equation (5.27)], [32, Section 4.5] to obtain the error convergence rate $n^{-\alpha}(\log n)^{\alpha d}$ or alternatively resort to higher-order digital nets [15] to achieve the exactly optimal rate of $n^{-\alpha}(\log n)^{(d-1)/2}$.
@@ 7. Concluding remarks
where $h>0$. These formulas are known to exhibit fast rates of convergence for functions analytic in certain regions of the complex plane [39]. Even the optimality of the double exponential formula is shown in [35] for analytic functions that decay as $\exp(-C|x|^{\rho})$, $\rho\geq 1$, on the real line. We also refer to [14] for recent further results for analytic functions.
@@
$\psi_{\textup{SE}}(x)=\tanh(x)\quad\text{ and }\quad\psi_{\textup{DE}}(x)=\tanh\bigg{(}\frac{\pi}{2}\sinh(x)\bigg{)}.$

### v2

@@ Abstract.
This work was supported by the Research Council of Finland (decisions 338567, 348503, 353081, 359181 and 359183).
@@ 1. Introduction
Let us briefly review recent literature on numerical integration and function approximation for finitely smooth functions over unbounded domains. When the weight is a Gaussian probability density, the corresponding weighted Sobolev spaces have been considered by numerous authors; see, e.g., [7, 20, 6, 12, 13] and the references therein. Freud weights, one possible generalization of Gaussian weights, have also been studied in our general context [10, 9], as have unweighted Sobolev spaces with a certain decay condition [31]. All aforementioned articles use the same weight for defining the Sobolev space for the target function and measuring the error in the considered approximation, which is an assumption that can be dropped: [29] employs unweighted Sobolev spaces for the target function but presents weighted $L^{2}$ and $L^{\infty}$ error estimates, [27] studies the relation between the two weights in a general framework, and univariate optimal algorithms and their convergence rates are considered in [24, 21]. Table 1 summarizes the settings of those aforementioned papers whose foci are close to ours.
@@
Concerning results in related settings, there have been studies on periodization strategies which transform a non-periodic target function defined on a finite closed interval into a periodic one. Combined with a subsequent use of the trapezoidal rule, one can attain a faster convergence rate. We refer to [36, Section 1] for an overview. However, we emphasize that for such studies on finitely smooth functions, the smoothness of the target function is often required as an input for the algorithm to attain the desired rate of convergence, whereas our method does not require this information and still achieves the optimal rate of convergence automatically. The multivariate counterpart of periodization strategy has been studied in the context of quasi-Monte Carlo methods, more precisely, lattice rules (see, e.g., [36, 26]). For instance, [16] shows that tent-transformed lattice rules can achieve second-order convergence in an appropriate function space setting, without any dimension dependence.
@@
The rest of this paper is organized as follows. In Section 2, we introduce and prove necessary concepts and useful lemmas related to monotonic Schwartz weights, Möbius transformations, and Sobolev spaces. Section 3 presents our main result for numerical integration, i.e., that the Möbius-transformed trapezoidal rule achieves the optimal rate of convergence. Section 4 extends this result to a randomized setting. In Section 5, we consider a problem of function approximation and prove the optimality of an algorithm that is based on combining the Möbius-transformed trapezoidal rule with trigonometric interpolation. Section 6 briefly considers a multidimensional extension of our method. Finally, Section 7 presents the concluding remarks. Induction proofs for a few technical results that are utilized in our analysis are collected in Appendix A.
@@ 2.1. Monotonic Schwartz weights
$x^{\alpha}\frac{|\omega^{(\beta)}(x)|^{r}}{\omega(x)}\leq(C^{\prime\prime\prime}_{\beta,m,\omega})^{r}x^{\alpha}\omega(x)^{mr/(m+\beta)-1}\longrightarrow 0$
@@
As (2.7) immediately follows from the definition of $\mathcal{S}$ for $\beta=0$, we may assume that $\beta\in\mathbb{N}$. A straightforward induction argument, presented in Appendix A, shows that for any $s>0$, the derivative $(\omega^{s})^{(\beta)}$ is a finite linear combination of terms of the form
@@ 2.3. Sobolev spaces
for some $1<q<\infty$ and $\alpha\in\mathbb{N}$, with $f^{(\tau)}$ denoting the $\tau$th weak derivative of $f$. The fact that $W^{\alpha,q}_{\rho}(\mathbb{R})$ is a Banach space (or a Hilbert space for $q=2$) for $\rho\in\mathcal{S}_{+}^{\rm mon}$ is a consequence of [23, Theorem 1.1 & Remark 4.10]. In what follows, we always consider the continuous representative of any given $f\in W^{\alpha,q}_{\rho}(\mathbb{R})$, which is possible for $\alpha\in\mathbb{N}$ due to the Sobolev embedding theorem and the inclusion $W^{\alpha,q}_{\rho}(\mathbb{R})\subset W^{1,q}_{\rm loc}(\mathbb{R})$. In Section 3, we specifically consider the case $q=2$ for numerical integration, even though the results also hold for any $q\in(1,\infty)$.
@@
The point evaluations at 00 and $2\pi$ in (2.3) are well-defined due to the trace theorem [1, Theorem 7.53] or the continuous embedding $W^{\alpha,q}(0,2\pi)\hookrightarrow C^{\alpha-1,\lambda}([0,2\pi])$ for Hölder indices $0<\lambda\leq 1-1/q$ [1, Part II in Theorem 5.4], demonstrating also that $W^{\alpha,q}_{\rm per}(0,2\pi)$ is a Banach space (or a Hilbert space for a $q=2$) as a closed subspace of $W^{\alpha,q}(0,2\pi)$. The closure of the smooth compactly supported test functions $C^{\infty}_{c}(0,2\pi)$ in the topology of $W^{\alpha,q}(0,2\pi)$ is denoted by $\mathring{W}^{\alpha,q}(0,2\pi)\subset W^{\alpha,q}_{\rm per}(0,2\pi)$. Note that in (2.14) the weak derivative $f^{(\tau)}$ is defined via partial integration of periodic functions on $\mathbb{T}$ (see, e.g., [33, Section 5.2]), whereas in (2.3), $f^{(\tau)}$ denotes the standard weak derivative on $(0,2\pi)\subset\mathbb{R}$, with $C^{\infty}_{c}(0,2\pi)$ serving as the space of test functions.
@@
For $1<q<\infty$ and $\alpha\in\mathbb{N}$, the spaces $W^{\alpha,q}(\mathbb{T})$ and $W^{\alpha,q}_{\rm per}(0,2\pi)$ can be identified via a bijective linear isometry $T:W^{\alpha,q}(\mathbb{T})\to W^{\alpha,q}_{\rm per}(0,2\pi)$.
@@
Let $\varphi:\mathbb{R}\ni\theta\mapsto(\cos\theta,\sin\theta)$ be the standard $2\pi$-periodic parametrization with respect to a polar angle $\theta$ for the embedding of $\mathbb{T}$ into $\mathbb{R}^{2}$. As $|\varphi^{\prime}|\equiv 1$, it is straightforward to conclude that the linear mapping $T:g\mapsto g\circ\varphi$ defines a bijective isometry from $W^{\alpha,q}(\mathbb{T})$ to
@@
for all $2\pi$-periodic $\psi\in C^{\infty}(\mathbb{R})$ (cf. [33, (5.5), (5.13) & Exercise 5.3.2]). Hence, the assertion follows if we prove that $\widetilde{W}^{\alpha,q}_{\rm per}(0,2\pi)=W^{\alpha,q}_{\rm per}(0,2\pi)$. This boils down to showing that the weak derivatives of $f\in W^{\alpha,q}_{\rm per}(0,2\pi)$ satisfy (2.16) and that those of $f\in\widetilde{W}^{\alpha,q}_{\rm per}(0,2\pi)$ are compatible with the conditions on point evaluations at 00 and $2\pi$ in (2.3).
@@
for all $2\pi$-periodic $\psi\in C^{\infty}(\mathbb{R})$, which could be proved by, e.g., approximating $f$ with smooth functions, integrating by parts and employing the continuous embedding $W^{\alpha,q}(0,2\pi)\hookrightarrow C^{\alpha-1,\lambda}([0,2\pi])$ for $0<\lambda\leq 1-1/q$ (cf. [1, Theorem 3.18 & Part II of Theorem 5.4]). The claim now follows by choosing in turns $f\in W^{\alpha,q}_{\rm per}(0,2\pi)$ and $f\in\widetilde{W}^{\alpha,q}_{\rm per}(0,2\pi)$ in (2.3) and comparing with (2.3) and (2.16). ∎
@@ 3. Möbius-Transformed Trapezoidal Rule
Let $\alpha\in\mathbb{N}$, $\rho\in\mathcal{S}_{+}^{\rm mon}$ and $f\in W^{\alpha,2}_{\rho}(\mathbb{R})$. Then it holds that
(3.2) $\big{|}I_{\rho}(f)-Q_{\rho,n}(f)\big{|}\leq C_{\rho,\alpha}n^{-\alpha}\|f\|_{W^{\alpha,2}_{\rho}(\mathbb{R})},$
@@
The proof relies on showing that $g=((f\rho)\circ\phi)\phi^{\prime}$ belongs to the $2\pi$-periodic $L^{2}$-based Sobolev space $W^{\alpha,2}_{\rm per}(0,2\pi)\cong W^{\alpha,2}(\mathbb{T})$ if $f\in W^{\alpha,2}_{\rho}(\mathbb{R})$ for $\alpha\in\mathbb{N}$. Indeed, after proving this, a classical result for the trapezoidal rule on periodic Sobolev spaces guarantees the convergence rate of order $n^{-\alpha}$; see, e.g.,  [2, Proposition 7.5.6] and [47, Theorem 2.4.1].
@@
Let us start with two auxiliary results that are proved by straightforward induction arguments in Appendix A. First, the weak derivative $g^{(\tau)}$, $\tau\in\mathbb{N}_{0}$, on the open interval $(0,2\pi)$ is a finite linear combination of terms of the form
@@
Figure 2. Absolute integration error for the logistic weight and $f(x)=|x|^{p}$, which corresponds to $I_{\rho}(f)=-2\,p!\,{\rm Li}_{p}(-1)$ where ${\rm Li}$ denotes the polylogarithm [19, Corollary 4.2]. The blue line shows the error for the Gauss–Logistic quadrature from [11]. The Möbius-transformed trapezoidal rule (green) exhibits much faster convergence.
@@ 4. Randomized trapezoidal rule
The convergence rate $n^{-\alpha-1/2}$ of the RMSE for the randomized quadrature rule (4.1) is a direct consequence of Lemma 3.2 combined with the classical result by Bakhvalov, [3] or [22, Theorem 11] where multidimensional settings are considered. Since our setting is one-dimensional, we refer to [13, p. 1670] for a better bound on the RMSE for $2T$-periodic functions. For completeness, we formally state this result as follows:
@@ 5. $L^{p}_{\rho}$ approximation
is as small as possible for $1\leq p<q<\infty$.
@@
To deduce the convergence rate of Theorem 5.2, we generalize Lemma 3.2 to show that the transformed target function $g_{p}$ belongs to the periodic $L^{q}$-based Sobolev space $W^{\alpha,q}_{\rm per}(0,2\pi)\cong W^{\alpha,q}(\mathbb{T})$ for $q>p$. The proof of Theorem 5.2 then follows from a trigonometric interpolation result by Temlyakov [47, Theorem 2.7].
@@
A straightforward induction argument, presented in Appendix A, demonstrates that the weak derivative $g_{p}^{(\tau)}$, $\tau\in\mathbb{N}$, on the open interval $(0,2\pi)$ is a finite linear combination of terms of the form
@@
We follow the line of reasoning in [6, Proof of Theorem 3.3]. To this end, let $f\in W^{\alpha,q}_{\rm per}(0,1)$ and denote its $1$-periodic extension onto the real line by the same symbol. An argument similar to that in the proof of Proposition 2.3 demonstrates that the extension $f$ belongs to $W^{\alpha,q}_{\rm loc}(\mathbb{R})$. For any $N\in\mathbb{N}$,
@@
The claim then follows from a lower bound for the approximation of periodic functions [47, Theorem 2.1.1], presented originally in [28]. ∎
@@ 6. Multivariate extension by componentwise transforms
Let $\alpha\in\mathbb{N}$, $\rho(\boldsymbol{x})=\prod_{k=1}^{d}\rho_{k}(x_{k})$ with $\rho_{k}\in\mathcal{S}_{+}^{\rm mon}$, $f\in W^{\alpha,2}_{\rho,\otimes}(\mathbb{R}^{d})$, and
@@
As the modified integrand function in (6.1) is in the periodic Sobolev space $W^{\alpha,2}_{\otimes}(\mathbb{T}^{d})$ that is norm-equivalent to the Sobolev space of dominating mixed smoothness $W^{\alpha,2}_{\rm mix}(\mathbb{T}^{d})$, one can employ in (6.1) good rank-$1$ lattice points [30, Equation (5.27)], [37, Section 4.5] to obtain the error convergence rate $n^{-\alpha}(\log n)^{\alpha d}$ or alternatively resort to higher-order digital nets [15] to achieve the exactly optimal rate of $n^{-\alpha}(\log n)^{(d-1)/2}$.
@@ 7. Concluding remarks
where $h>0$. These formulas are known to exhibit fast rates of convergence for functions analytic in certain regions of the complex plane [44]. Even the optimality of the double exponential formula is shown in [40] for analytic functions. We also refer to [14] for recent further results for analytic functions.
@@
$\psi_{\textup{SE}}(x)=\tanh\bigg{(}\frac{x}{2}\bigg{)}\quad\text{and }\quad\psi_{\textup{DE}}(x)=\tanh\bigg{(}\frac{\pi}{2}\sinh(x)\bigg{)}.$
@@
Acknowledgment
We thank Yoshihito Kazashi for his valuable comments.
Appendix A Induction proofs
The purpose of this appendix is to provide induction proofs for four technical results used in the above analysis, namely the representations (2.8), (3.5), (3.6) and (5.6).
Let us first prove (2.8), i.e., that for any $\beta\in\mathbb{N}$, the derivative $(\omega^{s})^{(\beta)}$ is a finite linear combination of terms of the form
(A.1) $\omega^{s}\prod_{j=1}^{\gamma}\frac{\omega^{(\tau_{j})}}{\omega},\qquad\text{with}\ \ \gamma\leq\beta\ \ \text{and}\ \ \sum_{j=1}^{\gamma}\tau_{j}=\beta.$
First of all, for $\beta=1$,
$(\omega^{s})^{\prime}=s\omega^{s}\frac{\omega^{\prime}}{\omega},$
which is of the required form with $\gamma=1$ and $\tau_{1}=1$. Assume then that the claim holds for an arbitrary but fixed $\beta\in\mathbb{N}$. The proof is completed by showing that the derivative of a term of the form (A.1) is a linear combination of terms that satisfy the same conditions with $\beta$ replaced by $\beta+1$:
$\bigg{(}\omega^{s}\prod_{j=1}^{\gamma}\frac{\omega^{(\tau_{j})}}{\omega}\bigg{)}^{\prime}$ $=s\omega^{s}\bigg{(}\frac{\omega^{\prime}}{\omega}\prod_{j=1}^{\gamma}\frac{\omega^{(\tau_{j})}}{\omega}\bigg{)}+\sum_{k=1}^{\gamma}\omega^{s}\prod_{j=1}^{\gamma}\frac{\omega^{(\tau_{j}+\delta_{jk})}}{\omega}-\gamma\omega^{s}\bigg{(}\frac{\omega^{\prime}}{\omega}\prod_{j=1}^{\gamma}\frac{\omega^{(\tau_{j})}}{\omega}\bigg{)}$
$=(s-\gamma)\omega^{s}\bigg{(}\frac{\omega^{\prime}}{\omega}\prod_{j=1}^{\gamma}\frac{\omega^{(\tau_{j})}}{\omega}\bigg{)}+\sum_{k=1}^{\gamma}\omega^{s}\prod_{j=1}^{\gamma}\frac{\omega^{(\tau_{j}+\delta_{jk})}}{\omega},$
where $\delta_{jk}$ is the Kronecker delta. As all summands on the right-hand side are of the required form, with the first one having $\gamma+1\leq\beta+1$ terms in its product and the others $\gamma$ terms, the assertion follows.
We then prove (3.5), i.e., that for $g=((f\rho)\circ\phi)\phi^{\prime}$ the derivative $g^{(\tau)}$, $\tau\in\mathbb{N}_{0}$, is a finite linear combination of terms of the form
(A.2) $\big{(}(f^{(\tau_{1})}\rho^{(\tau_{2})})\circ\phi\big{)}\prod_{j=1}^{\tau_{1}+\tau_{2}+1}\phi^{(\tau_{3,j})},\quad\text{with}\ \ \tau_{1}+\tau_{2}\leq\tau\ \ \text{and}\ \sum_{j=1}^{\tau_{1}+\tau_{2}+1}\tau_{3,j}=\tau+1.$
The case $\tau=0$ obviously holds with $\tau_{1},\tau_{2}=0$ and $\tau_{3,1}=1$. Assume then that the claim is true for an arbitrary but fixed $\tau\in\mathbb{N}_{0}$. The proof is completed by showing that the derivative of a term of the form (A.2) is a linear combination of terms that satisfy the same conditions with $\tau$ replaced by $\tau+1$:
$\bigg{(}\big{(}(f^{(\tau_{1})}\rho^{(\tau_{2})})\circ\phi\big{)}\prod_{j=1}^{\tau_{1}+\tau_{2}+1}$ $\phi^{(\tau_{3,j})}\bigg{)}^{\prime}=\sum_{k=1}^{\tau_{1}+\tau_{2}+1}\big{(}(f^{(\tau_{1})}\rho^{(\tau_{2})})\circ\phi\big{)}\prod_{j=1}^{\tau_{1}+\tau_{2}+1}\phi^{(\tau_{3,j}+\delta_{jk})}$
$+\big{(}(f^{(\tau_{1}+1)}\rho^{(\tau_{2})}+f^{(\tau_{1})}\rho^{(\tau_{2}+1)})\circ\phi\big{)}\bigg{(}\phi^{\prime}\prod_{j=1}^{\tau_{1}+\tau_{2}+1}\phi^{(\tau_{3,j})}\bigg{)}.$
As all summands on the right-hand side are of the required form, with either $\tau_{1}$ or $\tau_{2}$ increasing by one on the second line and neither of the two increasing on the first line when increasing the order of the derivative from $\tau$ to $\tau+1$, the assertion follows.
Then it is the turn of (5.6), i.e., we aim to prove that for $g_{p}=((f\,\rho^{1/p})\circ\phi)(\phi^{\prime})^{1/p}$ the derivative $g_{p}^{(\tau)}$, $\tau\in\mathbb{N}$, is a finite linear combination of terms of the form
(A.3) $\big{(}(f^{(\tau_{1})}(\rho^{1/p})^{(\tau_{2})})\circ\phi\big{)}(\phi^{\prime})^{1/p-\tau_{3}}\prod_{j=1}^{\tau_{1}+\tau_{2}+\tau_{3}}\phi^{(\tau_{4,j})},$
with
(A.4) $\tau_{1}+\tau_{2}+\tau_{3}\leq\tau\quad\text{and}\quad\sum_{j=1}^{\tau_{1}+\tau_{2}+\tau_{3}}\tau_{4,j}=\tau+\tau_{3}.$
To prove the case $\tau=1$, write
$\Big{(}\big{(}(f\,\rho^{1/p})\circ\phi\big{)}(\phi^{\prime})^{1/p}\Big{)}^{\prime}$ $=\big{(}(f^{\prime}\rho^{1/p}+f(\rho^{1/p})^{\prime})\circ\phi\big{)}(\phi^{\prime})^{1/p}\phi^{\prime}$
$\qquad+\frac{1}{p}\big{(}(f\,\rho^{1/p})\circ\phi\big{)}(\phi^{\prime})^{1/p-1}\phi^{(2)},$
where all three terms are of the required form with $(\tau_{1},\tau_{2},\tau_{3},\tau_{4,1})=(1,0,0,1)$, $(0,1,0,1)$ and $(0,0,1,2)$, respectively. Assume then that the claim is true for an arbitrary but fixed $\tau\in\mathbb{N}$. The proof is completed by showing that the derivative of a term of the form (A.3) is a linear combination of terms that satisfy the same conditions (A.3)-(A.4) with $\tau$ replaced by $\tau+1$:
[…]
$=\big{(}\big{(}f^{(\tau_{1}+1)}(\rho^{1/p})^{(\tau_{2})}+f^{(\tau_{1})}(\rho^{1/p})^{(\tau_{2}+1)}\big{)}\circ\phi\big{)}(\phi^{\prime})^{1/p-\tau_{3}}\bigg{(}\phi^{\prime}\prod_{j=1}^{\tau_{1}+\tau_{2}+\tau_{3}}\phi^{(\tau_{4,j})}\bigg{)}$
$\qquad+(1/p-\tau_{3})\big{(}(f^{(\tau_{1})}(\rho^{1/p})^{(\tau_{2})})\circ\phi\big{)}(\phi^{\prime})^{1/p-(\tau_{3}+1)}\bigg{(}\phi^{(2)}\prod_{j=1}^{\tau_{1}+\tau_{2}+\tau_{3}}\phi^{(\tau_{4,j})}\bigg{)}$
$\qquad+\sum_{k=1}^{\tau_{1}+\tau_{2}+\tau_{3}}\big{(}(f^{(\tau_{1})}(\rho^{1/p})^{(\tau_{2})})\circ\phi\big{)}(\phi^{\prime})^{1/p-\tau_{3}}\prod_{j=1}^{\tau_{1}+\tau_{2}+\tau_{3}}\phi^{(\tau_{4,j}+\delta_{jk})}.$
As all summands on the right-hand side are of the required form, with either $\tau_{1}$ or $\tau_{2}$ increasing by one on the first line, $\tau_{3}$ increasing by one on the second line and none of the three indices increasing on the final line when increasing the order of the derivative from $\tau$ to $\tau+1$, the assertion follows.
This appendix is completed by proving (3.6), i.e., that the derivatives of $\phi(\theta)=-c\cot(\theta/2)$ satisfy
(A.5) $\phi^{(\tau)}(\theta)=\frac{\psi_{\tau}(\theta)}{\sin^{\tau+1}(\theta/2)},\qquad\tau\in\mathbb{N}_{0},$
with $\psi_{\tau}(\theta)\in C^{\infty}(\mathbb{R})$ being a bounded finite linear combination of products of trigonometric functions. By definition, the claim holds for $\tau=0$. Assume that $\phi^{(\tau)}$ is of the form (A.5) for an arbitrary but fixed $\tau\in\mathbb{N}_{0}$ and let us differentiate:
$\phi^{(\tau+1)}(\theta)=\frac{\psi_{\tau}^{\prime}(\theta)\sin(\theta/2)-\tfrac{\tau+1}{2}\psi_{\tau}(\theta)\cos(\theta/2)}{\sin^{\tau+2}(\theta/2)},$
which immediately proves the claim.